# Theory of Mind abilities of Large Language Models in Human-Robot Interaction : An Illusion?

### 10 Jan 2024 v1

@@ Abstract.
Large Language Models have shown exceptional generative abilities in various natural language and generation tasks. However, possible anthropomorphization and leniency towards failure cases have propelled discussions on emergent abilities of Large Language Models especially on Theory of Mind (ToM) abilities in Large Language Models. While several false-belief tests exists to verify the ability to infer and maintain mental models of another entity, we study a special application of ToM abilities that has higher stakes and possibly irreversible consequences : Human Robot Interaction. In this work, we explore the task of Perceived Behavior Recognition, where a robot employs a Large Language Model (LLM) to assess the robot’s generated behavior in a manner similar to human observer. We focus on four behavior types, namely - explicable, legible, predictable, and obfuscatory behavior which have been extensively used to synthesize interpretable robot behaviors. The LLMs goal is, therefore to be a human proxy to the agent, and to answer how a certain agent behavior would be perceived by the human in the loop, for example ”Given a robot’s behavior X, would the human observer find it explicable?”. We conduct a human subject study to verify that the users are able to correctly answer such a question in the curated situations (robot setting and plan) across five domains. A first analysis of the belief test yields extremely positive results inflating ones expectations of LLMs possessing ToM abilities. We then propose and perform a suite of perturbation tests which breaks this illusion, i.e. Inconsistent Belief, Uninformative Context and Conviction Test. We conclude that, the high score of LLMs on vanilla prompts showcases its potential use in HRI settings, however to possess ToM demands invariance to trivial or irrelevant perturbations in the context which LLMs lack. We report our results on GPT-4 and GPT-3.5-turbo.
copyright: acmcopyrightjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYbooktitle: Woodstock ’18: ACM Symposium on Neural Gaze Detection, June 03–05, 2018, Woodstock, NYprice: 15.00isbn: 978-1-4503-XXXX-X/18/06
@@
Figure 1. Fetch executes the maneuver to pick the block and takes a step (towards robot’s) left indicating its choice of goal as Loc2 (instead of the alternative Loc3) thereby being legible. Left : shows the Fetch robot executing the maneuver. Center : the Fetch robot has a label on its head stating ”Not a Legible robot” but in a small font so the human observer is unable to read. Right : the human observer has a blurred view of the Fetch robot making them unable to comment on whether the robot is being legible or not. LLMs as human proxy expects the LLMs to ”think what the ’human would think of the robot’s behavior’”
@@ 1. Introduction
As Artificial Intelligence (AI) progresses, the development of the next generation of AI agents require the ability to interact with humans in human-like manner, processes and behaviors. A vital component of such interaction is the Theory of Mind (ToM), which involves attributing mental states – such as beliefs, intentions, desires, and emotions – to oneself and others, and to understand that these mental states may differ from one’s own. Theory of Mind has extensive roots in Human-Human interaction and has motivated several critical studies in pursuit of social intelligence (Astington and Jenkins, 1995; Astington and Baird, 2005; Frith and Frith, 2005; Wellman, 1992). Moreover, ToM is the corner stone enabling effective communication, collaboration, or deception and becomes a pre-requisite for most of human-agent interaction (Scassellati, 2001; Vinanzi et al., 2019; Hiatt et al., 2011; Devin and Alami, 2016; Lee et al., 2019). Everyday interactions which are second nature to human conversations like our ability to empathize with a character in a movie or understand social humor is in part due to our ability to perform theory of mind and comes naturally to humans (Sap et al., 2022).
@@
Theory of Mind becomes all the more important in Human-Robot Interaction to facilitate improved behavior synthesis (Chakraborti et al., 2018; Zhang et al., 2017; Chakraborti et al., 2017, 2015; Sreedharan et al., 2017, 2022; Kulkarni et al., 2019b, a, 2021) and engender Human-Robot trust (Zahedi et al., 2021, 2023, 2022b, 2022a; Esterwood and Robert, 2023). Typically, prior works have either assumed a human mental model (Chakraborti et al., 2018) or learned it through interaction (Zhang et al., 2017; Verma and Metcalf, 2022). Recent developments in generative AI, specifically Large Language Models may have opened a new way to access approximate human mental model to enable robots achieve their HRI objectives in better ways. LLMs are known to have excellent generative abilities. Initial research hinted at emergent LLM’s Reasoning, Planning & Theory of Mind abilities (Song et al., 2022; Huang et al., 2022; Liu et al., 2023a; Bubeck et al., 2023a), however, recent works have refuted such claims (Ullman, 2023a; Valmeekam et al., 2023; Shapira et al., 2023). While LLMs may not be sound reasoners or perform ToM as humans do, their performance on existing benchmarks for such tasks certainly exceeds chance behavior and it may find significant use as a Human-Proxy available to the robot. If LLMs were good at Theory of Mind then, as in Fig. 2, a robot can leverage the LLM to perform an internal critique while synthesizing behavior with a human-in-the-loop. The internal critique essentially would answer whether the planned robot behavior would be useful for the human-in-the-loop. For instance, if the human prefers explicable behavior (Zhang et al., 2017), the robot can query the LLM to check whether a planned behavior would indeed be explicable to the human.
@@
While past works checking for emergent abilities of LLM (Song et al., 2022; Huang et al., 2022; Liu et al., 2023a; Bubeck et al., 2023a), even ToM abilities (Kosinski, 2023a) exist, we posit that there is a need for investigating scenarios specific to HRI as real-world robots function in a non-ergodic world with irreversible consequences and much higher stakes. To this end, we investigate PROBE task (Perceived RObot Bhavior rEcognition) as a means to understand theory of mind abilities of Large Language Models in the context of Human-Robot Interaction. For example, the ”Fetch” robot that is navigating to a location with a human observer in the loop has to tuck its arms and crouch before moving [^1] Anonymous link to YouTube Playlist : https://www.youtube.com/playlist?list=PLRlaEr5LYlDBFD_1_Xz9_IblPzw1O1Y_X. The lay-observer unfamiliar with the underlying dynamics of the robot may find the actions ”tuck & crouch” unnecessary and inexplicable. In our investigation, the robot employs an LLM to ask the test question ”Given behavior X, would the human observer find such a plan explicable?”.
@@
We study four robot behaviors which have been extensively discussed in HRI and mental modelling, namely - explicable, legible, predictable and obfuscatory behavior (referred to as “behavior types”). We leverage the rich literature along these behavior types and borrow five domains used in prior works and valuable to the HRI community, namely - Fetch, Passage Gridworld, Environment Design, Urban Search and Rescue, and Package Delivery. For each domain we construct four distinct situations corresponding to each behavior type. The robot constructs a prompt using relevant context like description of the domain, robot capabilities & goals, and a definition of the behavior type. The LLM is asked a binary question (Q1) on whether the human observer would find the agent plan to be explicable (or the corresponding behavior type) and a multi choice question (Q2). In our evaluation, upon correct answer to (Q1), the LLM is tasked with choosing one explanation from a set of options, explaining why the plan conforms to the predicted behavior type (Q2). As a baseline, we first test LLM on vanilla prompts : success on which is indicative of potential ToM abilities but may also inflate one’s expectations of LLMs. Second, we propose three tests of which two are perturbation tests as Inconsistent Belief and Uninformative Context, and a Conviction Test. Our second suite of tests elicit important conclusions on ToM abilities of LLMs and essentially break the supposed illusion.
@@
  1. (1)
    We conduct a first study of Theory of Mind abilities of Large Language Models in the context of interpretable behavior synthesis and Human-Robot Interaction.
  2. (2)
    We investigate along five domains and four behavior types which have been well motivated with respect to the need for mental modeling in HRI and curate a total of 20 situations.
  3. (3)
    We perform a user subject study to validate our curated situations as lay users are able to correctly answer ToM queries.
  4. (4)
    We propose second-order tests including two perturbation tests and the Conviction test which highlights that any ToM abilities of LLMs are illusionary.
  5. (5)
    We perform a user subject study to validate that humans observers observing a real robot (Fetch) will be unaffected by perturbations while answering ToM queries.
Figure 2. An LLM being used as a Human-Proxy by a Robot as an internal-critique for behavior synthesis.
Figure 3. Interface used for our User Study: Example showing the three questions for Legibility in Fetch Domain.
@@ 2. Related Work
Large Language Models : Large Language Models (Zhao et al., 2023) are pre-trained generative AI models that can take a piece of text as an input and generate text as an output. Typically, the input space is unconstrained, meaning a finite yet large set of tokens / text is a valid input for the GPT like Large Language Model. While this allows for arbitrary tasks to be framed as a linguistic query for the LLM, it is unclear whether data-driven next-word prediction style of LLMs (OpenAI, 2023; Workshop et al., 2023; Anil et al., 2023; Driess et al., 2023) training is reason enough to believe that LLMs would perform well on said arbitrary tasks. Over the past years several variants and versions of Large Language Models have been released such as the GPT family (OpenAI, 2023), LLAMA family (Touvron et al., 2023), PALM family (Driess et al., 2023; Anil et al., 2023), BLOOM (Workshop et al., 2023) and several others (Zhao et al., 2023). In this work we conduct our study with GPT-4 (OpenAI, 2023) as it is one of the strongest LLMs in terms of NLP task performance.
@@
LLMs, Theory of Mind and Emergent abilities :
LLMs have amassed a large popularity with public and the research community with its exceptional abilities in NLP tasks. With an AI as impressive as an LLM, researchers have explored the possibilities of emergent abilities of these LLMs. While this is an active area of research, of interest to the scope of this work are LLMs abilities in reasoning (Kojima et al., 2022; Wang et al., 2022; Wei et al., 2022), deception abilities (Hagendorff, 2023), Theory of Mind abilities (Kosinski, 2023b) and other emergent behaviors that are similar to human behaviors (Sap et al., 2019; Rashkin et al., 2018; Suresh and Ong, 2021; Yang et al., 2019; Bubeck et al., 2023b). While these works highlight how LLMs possess these abilities, more recent works have refuted these claims (Valmeekam et al., 2023; Ullman, 2023a; Borji, 2023; Kocoń et al., 2023; Shi et al., 2023; McCoy et al., 2023) citing that LLMs are extremely brittle and heavilty depend on the prompt fed to it. (Valmeekam et al., 2023) suggests, systems that reason do not have to depend upon such syntactic alterations to the input like Chain of Thought Prompts, Tree of Thoughts or other prompt engineering strategies (Liu et al., 2023b). While ToM has extensive prior works on specialized robot agents (Marsella et al., 2004; Pynadath and Marsella, 2005; Baker et al., 2011) these works are not directly applicable to investigating the emergent ToM abilities of LLMs. We distinguish ourselves from past works through our novel construction of prompt situations which are sourced from the vast literature on behavior synthesis, the situations are well rooted for HRI scenarios, provide a thorough analysis via a large-scale user subject study and provide a case study with real Fetch robot in HRI setting.
@@ 3. Preliminaries
A core focus of this work is to analyze the Theory of Mind abilities of Large Language models in Human-AI interaction scenario. As motivated previously, we identify that Human-Aware Planning is an import part of HRI that attempts to synthesize agent behaviors by taking human mental model into account. (Chakraborti et al., 2018; Sreedharan et al., 2022) provides a good overview of the literature and formal specifications of behavior types we consider in this work. (Chakraborti et al., 2018; Kulkarni et al., 2019a) papers reconciles the various interpretations of terminologies used in the literature and highlights important connections between related works. Subsequently, they provide formal definitions of the four behavior types (explicability, predictability, legibility and obfuscation) as follows :
@@ 4. Methodology
Research Question 1: How do lay users perform on ToM reasoning tasks in Human-Robot Interaction scenarios?
Research Question 2: How much does LLM performance align with that of: 1) oracle, & 2) lay users?
@@
As discussed in Sec. 6, the performance of LLMs on the vanilla prompts (where $P=\emptyset$) engenders an inflated belief that LLMs may possess ToM abilities. We then investigate the robustness of LLM responses with two perturbation cases and a Conviction test. Unlike how ToM agents (like humans) behave (Astington and Jenkins, 1995; Astington and Baird, 2005; Frith and Frith, 2005; Wellman, 1992; Ullman, 2023a)(see Sec. 7), we find that LLMs are extremely brittle to small changes to prompt and fail miserably. Finally, we present a case study with Fetch robot where a human observes the robot acting in the world along the three cases Vanilla prompt, Uninformative Context and Inconsistent Belief. For all our experiments we use the most recent version of GPT-4 (as of September 30, 2023) and GPT-3.5-turbo (as of November 25, 2023). Past works have justified their general superiority amongst the popular LLMs and within the GPT-X family (OpenAI, 2023) motivating their use as representative LLMs.

### 17 Jan 2024 v2

@@ Abstract.
Large Language Models (LLMs) have shown exceptional generative abilities in various natural language and generation tasks. However, possible anthropomorphization and leniency towards failure cases have propelled discussions on emergent abilities of LLMs especially on Theory of Mind (ToM) abilities in Large Language Models. While several false-belief tests exists to verify the ability to infer and maintain mental models of another entity, we study a special application of ToM abilities that has higher stakes and possibly irreversible consequences : Human Robot Interaction. In this work, we explore the task of Perceived Behavior Recognition, where a robot employs an LLM to assess the robot’s generated behavior in a manner similar to human observer. We focus on four behavior types, namely - explicable, legible, predictable, and obfuscatory behavior which have been extensively used to synthesize interpretable robot behaviors. The LLMs goal is, therefore to be a human proxy to the agent, and to answer how a certain agent behavior would be perceived by the human in the loop, for example ”Given a robot’s behavior X, would the human observer find it explicable?”. We conduct a human subject study to verify that the users are able to correctly answer such a question in the curated situations (robot setting and plan) across five domains. A first analysis of the belief test yields extremely positive results inflating ones expectations of LLMs possessing ToM abilities. We then propose and perform a suite of perturbation tests which breaks this illusion, i.e. Inconsistent Belief, Uninformative Context and Conviction Test. The high score of LLMs on vanilla prompts showcases its potential use in HRI settings, however to possess ToM demands invariance to trivial or irrelevant perturbations in the context which LLMs lack. We report our results on GPT-4 and GPT-3.5-turbo.
Theory of Mind; Large Language Models; Reasoning
journalyear: 2024copyright: acmlicensedconference: Companion of the 2024 ACM/IEEE International Conference on Human-Robot Interaction; March 11–14, 2024; Boulder, CO, USAbooktitle: Companion of the 2024 ACM/IEEE International Conference on Human-Robot Interaction (HRI ’24 Companion), March 11–14, 2024, Boulder, CO, USAdoi: 10.1145/3610978.3640767isbn: 979-8-4007-0323-2/24/03ccs: Computing methodologies Theory of mindccs: Computing methodologies Knowledge representation and reasoningccs: Computing methodologies Planning and scheduling
@@
Figure 1. Fetch executes the maneuver to pick the block and takes a step (towards robot’s) left indicating its choice of goal as Loc2 (instead of the alternative Loc3) thereby being legible. Left : Fetch executes the maneuver. Center : Fetch has a label on its head stating ”Not a Legible robot” in a small font which the human observer is unable to read. Right : Human has a blurred view of Fetch making them unable to comment on whether the robot is being legible. Robot expects the LLM (human proxy) to ”think what the ’human would think of the robot’s behavior’”
**footnotetext: These authors contributed equally to this work
(a) Using LLMs as a Human-Proxy
(b) User Study Interface
Figure 2. (a) An LLM being used as a Human-Proxy by a Robot as an internal-critique for behavior synthesis. (b) Interface used for our User Study: Example showing the three questions for Legibility in Fetch Domain.
@@ 1. Introduction
As Artificial Intelligence (AI) progresses, the development of the next generation of AI agents require the ability to interact with
@@
humans in human-like manner, processes and behaviors. A vital component of such interaction is the Theory of Mind (ToM), which involves attributing mental states – such as beliefs, intentions, desires, and emotions – to oneself and others, and to understand that these mental states may differ from one’s own. ToM has extensive roots in Human-Human interaction and has motivated several critical studies in pursuit of social intelligence (Astington and Jenkins, 1995; Astington and Baird, 2005; Frith and Frith, 2005; Wellman, 1992). Moreover, ToM is the corner stone enabling effective communication, collaboration, or deception and becomes a pre-requisite for most of human-agent interaction (Scassellati, 2001; Vinanzi et al., 2019; Hiatt et al., 2011; Devin and Alami, 2016; Lee et al., 2019). Everyday interactions which are second nature to human conversations like our ability to empathize with a character in a movie or understand social humor is in part due to our ability to perform theory of mind and comes naturally to humans (Sap et al., 2022).
@@
Theory of Mind becomes all the more important in Human-Robot Interaction to facilitate improved behavior synthesis (Chakraborti et al., 2018; Zhang et al., 2017; Chakraborti et al., 2017, 2015; Sreedharan et al., 2017, 2022; Kulkarni et al., 2019b, a, 2021) and engender Human-Robot trust (Zahedi et al., 2021, 2023, 2022b, 2022a; Esterwood and Robert, 2023). Prior works have either assumed a human mental model (Chakraborti et al., 2018) or learned it through interaction (Zhang et al., 2017; Verma and Metcalf, 2022). Recent developments in generative AI, specifically LLMs may have opened a new way to access approximate human mental model to enable robots achieve their HRI objectives in better ways through symbolic (Sreedharan et al., 2020) or other modalities (Guan et al., 2021; Verma and Metcalf, 2022; Verma et al., 2023). Initial research hinted at emergent LLM’s Reasoning, Planning & Theory of Mind abilities (Song et al., 2022; Huang et al., 2022; Liu et al., 2023a; Bubeck et al., 2023a), however, recent works have refuted such claims (Ullman, 2023a; Valmeekam et al., 2023; Shapira et al., 2023). While LLMs may not be sound reasoners or perform ToM as humans do, their performance on existing benchmarks for such tasks certainly exceeds chance behavior and it may find significant use as a Human-Proxy available to the robot. Sally-Anne test and its variations that have been extensively studied for testing ToM have likely been seen in the training data for these LLMs. We suspect that LLM’s approximate retrieval is be mistaken for reasoning abilities. If LLMs were indeed good at Theory of Mind, then, as in Fig. 1(a), a robot can leverage the LLM to perform an internal critique while synthesizing behavior with a human-in-the-loop. The internal critique essentially would answer whether the planned robot behavior would be useful for the human-in-the-loop. For instance, if the human prefers explicable behavior (Zhang et al., 2017), the robot can query the LLM to check whether a planned behavior would indeed be explicable to the human.
@@
While past works checking for emergent abilities of LLM (Song et al., 2022; Huang et al., 2022; Liu et al., 2023a; Bubeck et al., 2023a), even ToM abilities (Kosinski, 2023a) exist, we posit that there is a need for investigating scenarios specific to HRI as real-world robots function in a non-ergodic world with irreversible consequences and much higher stakes. To this end, we investigate PROBE task (Perceived RObot Bhavior rEcognition) as a means to understand theory of mind abilities of Large Language Models in the context of Human-Robot Interaction. For example, the Fetch robot that is navigating to a location with a human observer in the loop has to tuck its arms and crouch before moving [^1] Project Link & Supplementary : https://famishedrover.github.io/papers/llm-probe/index.html. The lay-observer unfamiliar with the underlying dynamics of the robot may find the actions ”tuck & crouch” unnecessary and inexplicable. In our investigation, the robot employs an LLM to ask the test question ”Given behavior X, would the human observer find such a plan explicable?”.
@@
We study four robot behaviors which have been extensively discussed in HRI and mental modelling, namely - explicable, legible, predictable and obfuscatory behavior (referred to as “behavior types”). We leverage the rich literature along these behavior types and borrow five domains used in prior works and valuable to the HRI community, namely - Fetch, Passage Gridworld, Environment Design, Urban Search and Rescue, and Package Delivery. For each domain we construct four distinct situations corresponding to each behavior type. We construct a prompt (can be automated) using relevant context like description of the domain, robot capabilities & goals, and a definition of the behavior type. The LLM is asked a binary question (Q1) on whether the human observer would find the agent plan to be explicable (or the corresponding behavior type) and a multi choice question (Q2). In our evaluation, upon correct answer to (Q1), the LLM is tasked with choosing one explanation from a set of options, explaining why the plan conforms to the predicted behavior type (Q2). As a baseline, we first test LLM on vanilla prompts : success on which is indicative of potential ToM abilities but may also inflate one’s expectations of LLMs. Second, we propose three tests of which two are perturbation tests as Inconsistent Belief and Uninformative Context, and a Conviction Test. Our second suite of tests elicit important conclusions on ToM abilities of LLMs and essentially break the supposed illusion.
@@
1. We conduct a first study of Theory of Mind abilities of Large Language Models in the context of interpretable behavior synthesis and Human-Robot Interaction.
2. We investigate along five domains and four behavior types which have been well motivated with respect to the need for mental modeling in HRI and curate a total of 20 situations.
3. We perform a user subject study to validate our curated situations as lay users are able to correctly answer ToM queries.
4. We propose second-order tests including two perturbation tests and the Conviction test which highlights that any ToM abilities of LLMs are illusionary.
5. We perform a user subject study to validate that humans observers observing a real robot (Fetch) will be unaffected by perturbations while answering ToM queries.
@@ 2. Related Work
Large Language Models : Large Language Models (Zhao et al., 2023) are pre-trained generative AI models that can take a piece of text as an input and generate text as an output. Typically, the input space is unconstrained, meaning a finite yet large set of tokens / text is a valid input for the GPT like Large Language Model. While this allows for arbitrary tasks to be framed as a linguistic query for the LLM, it is unclear whether data-driven next-word prediction style of LLMs (OpenAI, 2023b; Workshop et al., 2023; Anil et al., 2023; Driess et al., 2023) training is reason enough to believe that LLMs would perform well on said arbitrary tasks. Over the past years several variants and versions of Large Language Models have been released such as the GPT family (OpenAI, 2023b), LLAMA family (Touvron et al., 2023), PALM family (Driess et al., 2023; Anil et al., 2023), BLOOM (Workshop et al., 2023) and several others (Zhao et al., 2023). In this work we conduct our study with GPT-4 (OpenAI, 2023b) as it is one of the strongest LLMs in terms of NLP task performance.
@@
LLMs, Theory of Mind and Emergent abilities: LLMs have amassed a large popularity with public and the research community with its exceptional abilities in NLP tasks. With an AI as impressive as an LLM, researchers have explored the possibilities of emergent abilities of these LLMs. While this is an active area of research, of interest to the scope of this work are LLMs abilities in reasoning (Kojima et al., 2022; Wang et al., 2022; Wei et al., 2022), deception abilities (Hagendorff, 2023), Theory of Mind abilities (Kosinski, 2023b) and other emergent behaviors that are similar to human behaviors (Sap et al., 2019; Rashkin et al., 2018; Suresh and Ong, 2021; Yang et al., 2019; Bubeck et al., 2023b). While these works highlight how LLMs possess these abilities, more recent works have refuted these claims (Valmeekam et al., 2023; Ullman, 2023a; Borji, 2023; Kocoń et al., 2023; Shi et al., 2023; McCoy et al., 2023) citing that LLMs are extremely brittle and heavilty depend on the prompt fed to it. (Valmeekam et al., 2023) suggests, systems that reason do not have to depend upon such syntactic alterations to the input like Chain of Thought Prompts, Tree of Thoughts or other prompting strategies (Liu et al., 2023b). While ToM has extensive prior works on specialized robot agents (Marsella et al., 2004; Pynadath and Marsella, 2005; Baker et al., 2011) these works are not directly applicable to investigating the emergent ToM abilities of LLMs. We distinguish ourselves from past works through our novel construction of prompt situations which are sourced from the vast literature on behavior synthesis, the situations are well rooted for HRI scenarios, provide a thorough analysis via a large-scale user subject study and provide a case study with real Fetch robot in HRI.
@@ 3. Preliminaries
A core focus of this work is to analyze the Theory of Mind abilities of Large Language models in Human-AI interaction scenario. As motivated previously, we identify that Human-Aware Planning is an important part of HRI that attempts to synthesize agent behaviors by taking human mental model into account. (Chakraborti et al., 2018; Sreedharan et al., 2022) provides a good overview of the literature and formal specifications of behavior types we consider in this work. (Chakraborti et al., 2018; Kulkarni et al., 2019a) papers reconciles the various interpretations of terminologies used in the literature and highlights important connections between related works. Subsequently, they provide formal definitions of the four behavior types (explicability, predictability, legibility and obfuscation) as follows :
@@ 4. Methodology
Research Question 1: How do human subjects [^2] We refer to lay users as human subjects in the rest of the paper. perform on ToM reasoning tasks in Human-Robot Interaction scenarios?
Research Question 2: How much does LLM performance align with that of: 1) oracle, & 2) human subjects?
@@
As discussed in Sec. 6, the performance of LLMs on the vanilla prompts (where $P=\emptyset$) engenders an inflated belief that LLMs may possess ToM abilities. We then investigate the robustness of LLM responses with two perturbation cases and a Conviction test. Unlike how ToM agents (like humans) behave (Astington and Jenkins, 1995; Astington and Baird, 2005; Frith and Frith, 2005; Wellman, 1992; Ullman, 2023a)(see Sec. 7), we find that LLMs are extremely brittle to small changes to prompt and fail miserably. Finally, we present a case study with Fetch robot where a human observes the robot acting in the world along the three cases Vanilla prompt, Uninformative Context and Inconsistent Belief. For all our experiments we use the most recent version of GPT-4 (as of September 30, 2023) and GPT-3.5-turbo (as of November 25, 2023). Past works have justified their general superiority amongst the popular LLMs and within the GPT-X family (OpenAI, 2023b) motivating their use as representative LLMs.